OpenStudy (m_m13):

will medal and fan 3 questions

OpenStudy (anonymous):

?

OpenStudy (m_m13):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B. How fast did Car A travel? ________mph

OpenStudy (m_m13):

Gretchen and Ezia received equal scores on a test made up of multiple choice questions and an essay. Gretchen got 18 multiple choice questions correct and received 19 points for her essay. Ezia got 15 multiple choice questions correct and received 31 points for her essay. How many points was each multiple choice question worth? _______points

OpenStudy (m_m13):

In October, Greg and Thomas had the same amount of money in their savings accounts. In November, Greg deposited $120 into his account. Thomas increased the money in his account by 20%. When they compared their balances, they found that they were still equal. How much money did each person have in his account in October? $________

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

here is the answer for the first question:You have the formula, Velocity=distance/time hence we'll use that to solve for the velocity of both Cars A and B since Car A traveled for 2.4 hours, then we'll have Va = d/2.4 -----equation 1 since Car B traveled for 4 hours, then, Vb = d/4 -----equation 2 NOTE that the distance traveled by Cars A and B are equal getting the distance from equations 1 & 2 Va = d/2.4 d=2.4Va -------equation 3 Vb = d/4 d=4Vb -------equation 4 since distance of A and B are equal then we can equate equations 3&4 2.4Va=4Vb ------equation 5 The problem states that Car A traveled 22mph faster than car B. Therefore we'll have the equation as: Va=Vb+22 -------equation 6 substituting equation 6 to equation 5 we'll have 2.4Va=4Vb 2.4(Vb+22)=4Vb 2.4Vb +22(2.4)=4Vb 4Vb-2.4Vb=52.8 1.6Vb=52.8 Vb=33mph substituting Vb to equation 6 Va=Vb+22 Va=33+22 Va=55mph

OpenStudy (m_m13):

thanks

OpenStudy (anonymous):

you will have to set this up into two equations and then use substitution or elimination. so the first equation would be 18x+19=y can you figure out the second equation from here?

OpenStudy (anonymous):

this is for the second question